OpenStudy (tiffanymak1996):

find d/dx (integral[(x+y),(x-y)] sin(t^3) dt)

OpenStudy (tiffanymak1996):

\[\frac{ d }{ dx } \int\limits_{x+y}^{x-y} \sin(t ^{3}) dt\]

zepdrix (zepdrix):

This is a partial derivative again?

OpenStudy (tiffanymak1996):

yes...

OpenStudy (tiffanymak1996):

do i have to use chain rule for (x-y) and (x+y) or just leave it as sin (x-y)^3 -sin (x+y)^3?

zepdrix (zepdrix):

You `do` have to apply the chain rule in each case. But it looks the chain rule is just giving us 1. Sooooo, yah.

zepdrix (zepdrix):

If the limits on our integral had been something like (x-c) and (x+c), we would worry about applying the chain rule since the derivative of those limits is just 1. Same idea here since we're applying partials :)

zepdrix (zepdrix):

We wouldn't*

OpenStudy (tiffanymak1996):

thanks!